# Imagination Augmented Generation: Learning to Imagine Richer Context for Question Answering over Large Language Models

### v3

@@ Abstract
Retrieval-Augmented-Generation and Gener-ation-Augmented-Generation have been proposed to enhance the knowledge required for question answering over Large Language Models (LLMs). However, the former relies on external resources, and both require incorporating explicit documents into the context, which increases execution costs and susceptibility to noise data. Recent works indicate that LLMs have modeled rich knowledge, albeit not effectively triggered or awakened. Inspired by this, we propose a novel knowledge-augmented framework, Imagination-Augmented-Generation (IAG), which simulates the human capacity to compensate for knowledge deficits while answering questions solely through imagination, thereby awakening relevant knowledge in LLMs without relying on external resources. Guided by IAG, we propose an imagine richer context method for question answering (IMcQA). IMcQA consists of two modules: explicit imagination, which generates a short dummy document by learning from long context compression, and implicit imagination, which creates flexible adapters by distilling from a teacher model with a long context. Experimental results on three datasets demonstrate that IMcQA exhibits significant advantages in both open-domain and closed-book settings, as well as in out-of-distribution generalization111 Our code will be available at https://anonymous.4open.science/r/IMcQA.
@@ 1 Introduction
Figure 1: Compared with RAG and GAG , the proposed IAG eschews external resources, generates a shorter context (explicitly imagination) and creates flexible adapters (implicitly imagination) for each question.
@@
In recent years, to mitigate hallucinations in LLMs and improve performance in knowledge-intensive tasks such as QA, researchers have proposed numerous knowledge-augmented methods for LLMs. These methods primarily fall into two categories: Retrieval-Augmented-Generation (RAG) Guu et al. (2020) and Generation-Augmented-Generation (GAG) Kim et al. (2024). RAG (Top part of Figure 1) retrieves related documents from external resources (e.g., Wikipedia) and then inputs both the retrieved documents and the question into LLMs Izacard and Grave (2021). GAG (Middle part of Figure 1) leverages external LLMs such as ChatGPT Ouyang et al. (2022) to generate relevant documents Yu et al. (2023).
@@
However, these methods have the following disadvantages [^2] A more intuitive comparison can be seen in A.1.: 1) Dependence on external resources, RAG relies on external domain knowledge resources Ke et al. (2024), while GAG depends on a more powerful external LLM as a knowledge generator. This dependency hinders their broader application. 2) Increased execution costs, The required computing resources and inference time continue to increase as the number of documents increases Xu et al. (2023a). For example, the typical RAG method FiD Izacard and Grave (2021) needs to handle over 12K tokens for retrieving 100 documents, leading to more than a $100$-fold increase in prompt length and over $100^{2}$-fold increase in inference time Liu et al. (2023a). The GAG method Yu et al. (2023) requires additional financial costs (e.g., API calls). 3) Susceptibility to noise data, both methods are prone to the inclusion of irrelevant or noisy data Li et al. (2023); Shaier et al. (2024). For example, Gao et al. (2024) indicates that noise in the documents negatively affects the performance. Therefore, there is an urgent need to explore new knowledge-augmented methods.
@@
In fact, LLMs inherently contain rich knowledge and possess significant potential for resolving knowledge-intensive tasks Bhagavatula et al. (2020). Enhancing the performance of specific tasks can be achieved by better activating and awakening relevant knowledge or expanding memory capacity without relying on external resources. For example, simply repeating the question twice Xu et al. (2023b), just reviewing and consolidating knowledge by appending a straightforward prompt “As far as I know” Yao et al. (2023), and using visual-language models to imagine images Tang et al. (2023), all these approaches enhance the performance of LLMs on downstream tasks.
Therefore, inspired by the above phenomena and aiming to address the challenges faced by RAG and GAG, we propose a novel knowledge-augmented framework, Imagination-Augmented-Generation (IAG). IAG simulates the human capacity to compensate for knowledge deficits solely through imagination in QA. For resolving knowledge-intensive tasks, IAG utilizes solely LLMs to generate a shorter context and create a more flexible adapter (Bottom part of Figure 1). Through a knowledge distillation framework, IAG achieves knowledge enhancement through imagination. The fundamental idea is to permit student models, which are deficient in rich contextual information (e.g., without abundant retrieved documents), to imitate teacher models that are endowed with such information.
Within the framework of IAG, we introduce an imagine richer context method for question answering (IMcQA). To sufficiently utilize the inherent knowledge of LLMs, we design two main modules to activate the various potential knowledge modeling in LLMs and obtain a richer context. Specifically, the explicit imagination module first uses symbol distillation to obtain the compressed context and then guides the context generator to generate a short dummy document. Subsequently, within the knowledge distillation framework, the implicit imagination module utilizes the hypernetwork to generate adapters to awaken the question-aware ability of LLMs. Unlike the LoRA Hu et al. (2021) stores task knowledge and ability in modules, the hypernetwork learns to generate a question-aware knowledge module.
@@
We evaluate the proposed IMcQA on various LLMs, including T5 Roberts et al. (2020a) and Llama2 Touvron et al. (2023). The experimental results across three QA datasets indicate that the proposed method yields performance gains while reducing computational expenses and time. Notably, it even outperforms baseline methods that retrieve and generate knowledge under the same document settings. In conclusion, the contributions of this paper are summarized as follows:
@@
  • We propose a new knowledge augmentation framework IAG to fully leverage the LLMs’ intrinsic knowledge more efficiently without relying on external resources.
  • We propose a novel QA method IMcQA that employs two main modules (explicit imagination and implicit imagination) to utilize better the knowledge stored in the LLMs and awaken a richer context in QA.
@@
  • Experimental results indicate that the proposed method successfully activates the relevant internal knowledge of LLMs. IMcQA exhibits significant advantages in both open-domain and closed-book settings, as well as in out-of-distribution generalizations.
@@ 2 Related Work
Knowledge Enhancement has usually been adopted to alleviate the issue of insufficient knowledge in LLMs. There are two main methods: RAG Sun et al. (2019); Wang et al. (2024) and GAG Abdallah and Jatowt (2023). The typical RAG method FiD Izacard and Grave (2021) retrieves documents from Wikipedia to answer questions. LLMs serving as a knowledge base have been the focus of numerous studies that advocate the extraction of knowledge from such models (e.g., GPT-3). For instance, Yu et al. (2023) generates 10 documents for each question. However, RAG requires external resources, and both RAG and GAG need verbose long contexts. Recently, methods have been developed to enhance LLMs’ abilities by simulating human imagination of visual information using existing visual-language models Tang et al. (2023); Akter et al. (2024). We prefer self-imagination to augment knowledge and aim to leverage the parameterized knowledge within the models Xu et al. (2023b); Kazemnejad et al. (2023). Our proposed method not only eliminates the need for external resources but also improves the efficiency of activating internal knowledge within LLMs.
@@
Knowledge Distillation is a technique where a smaller model learns to mimic the predictions of a larger model, aiming to retain performance while reducing computational resources Hinton et al. (2015). Recent studies West et al. (2022) present symbolic knowledge distillation, a process that facilitates knowledge transfer from a teacher model via extracting training data to subsequently train a student model Wang et al. (2023b); Ranaldi and Freitas (2024). In this paper, the process of obtaining compressed context during the explicit imagination resembles a form of symbolic distillation. Regarding training, our emphasis lies in distilling the long-context modeling abilities of LLMs.
Hypernetworks is designed to reduce the number of parameters Ha et al. (2016), i.e., a small neural network generates parameters for another big neural network. Recent studies Phang et al. (2022); Ivison et al. (2023) have explored the enhancement of model performance in zero/few-shot settings through meta-learning involving hypernetworks. We utilize hypernetworks to acquire implicit imagine capabilities by dynamically generating LoRA for efficiency and generalization.
@@ 3 Method
In this section, we introduce the detailed method of IMcQA to activate LLMs’ intrinsic knowledge and obtain a richer context for QA. The fundamental premise underlying this method is that QA with a richer context yields greater performance (e.g., RAG with retrieved documents). Consequently, diverse methods are employed for questions lacking in richer contexts to activate knowledge within LLMs to replicate comparable effects to those achieved with richer contexts.
Specifically, IMcQA comprises two main modules. Explicit imagination with long context compression learns to imagine a short dummy document (§ 3.2). And implicit imagination with the hypernetwork models’ hidden knowledge that learns a shared knowledge feature projection across questions (§ 3.3). The hypernetwork is trained to generate lightweight LoRA modules, aiming to align the question and the internal knowledge. Besides, there is long context distillation in training, which learns the teacher’s rich representations to compensate for missing knowledge in imagination (§ 3.4).
@@ 3.1 Formulation
The formulation of our task follows RAG for QA Guu et al. (2020). Let $\mathcal{V}^{*}$ denote the infinite set, encompassing all potential strings over the tokens in vocabulary $\mathcal{V}$, and this includes the empty string. An instance within a QA dataset is defined as a triplet $(q,a,c)$ comprising question $q$, answer $a$, and context $c$, where $q,a,c\in\mathcal{V}^{*}$. Conventionally, the context $c$ is drawn from the knowledge corpus $\mathcal{Z}$, like Wikipedia, whereby $\mathcal{Z}\subset\mathcal{V}^{*}$. Additional background details are available in B.1.
Figure 2: Overview of IMcQA method. In the inference phase, for each question, the explicit imagination generates a short dummy document and the implicit imagination with hypernetwork creates a specific LoRA module. During training, there are two stages: the first stage is the pre-training of the context generator (§ 3.2), aiming at its ability to imagine a short dummy document based on the question, and the second stage is the hypernetwork fine-tuning (§ 3.3) using long context distillation (§ 3.4) to obtain a question-specific LoRA module.
3.2 Explicit Imagination with Context Generator
To obtain a context $c$ for each question, we utilize LLMs to generate a short dummy document, which can mitigate knowledge corpus error Lee et al. (2023) by considering potentially useful contexts. We employ the "teacher-student" framework. During the training phase, the teacher model, equipped with rich contextual information, generates short contexts through context compression. The student model, in turn, learns to automatically generate these short contexts by leveraging its own LLM capabilities.
As shown in the left part of Figure 2, to help LLMs fully utilize the knowledge and generate compressed text, we first pretrain the context generator on our collected question-compressed document pairs. By leveraging symbolic distillation, we employ the long context compression method LongLLMLingua Jiang et al. (2023) to condense a large corpus of retrieved documents. These compressed texts $c^{\prime}$ then serve as fine-tuning data alongside specific prompts $p_{q}$ (B.2) and question-answer pairs for the context generator [^3] We discuss the role of context generator in the A.2. $G_{\theta}$ ($\theta$ represents the model’s parameter), which guides the model to think about its knowledge and generate a short dummy document:
$d=G_{\theta}(p_{q}(q;c^{\prime}))$ (1)
where $d$ is the dummy document generated from the context generator. This process enables LLMs to conceive compressed knowledge that robustly parallels the question’s knowledge requirements.
3.3 Implicit Imagination with Hypernetwork
@@
We utilize the hypernetwork [^4] We conduct a detailed analysis of the reasons behind the hypernetwork in the A.3. to learn to generate a better QA model, which does not directly optimize the LoRA Hu et al. (2021) module but generates specific LoRA adapter weights using the inputs for QA (bottom part of Figure 2). This is akin to repeating the question in the prompt Xu et al. (2023b) and incorporating certain topical cues to stimulate the model’s recall of relevant questions Wang et al. (2023c). However, the distinction lies in the fact that they serve as wake-up features, whereas we are generating model parameters.
The hypernetwork architecture for generating LoRA weights is exhaustively outlined in Figure 3. Specifically, $D_{k}^{q}$ and $U_{k}^{q}$ represent the low-rank projections of layer $k$ correlated with the $\mathcal{Q}$, while $D_{k}^{v}$ and $U_{k}^{v}$ correspond to those associated with the $\mathcal{V}$. Hypernetwork represented as $g_{D}$ and $g_{U}$, takes $\textit{concat}(f,i_{k}^{\{q,v\}})$ as input, where $f$ is the feature vectors that use the model’s encoder to obtain and using whitening algorithm Su (2021) for dimensionality reduction, $i_{k}^{\{q,v\}}\in\{0,\ldots,2\times\#\text{blocks}\}$ signifies the positional embedding differentiating between layers and between $\mathcal{QV}$. Each hypernetwork is defined by weights $W_{d}$ and $W_{u}$ which represent the down and up projections respectively. Finally, the hypernetwork equations for $D^{\{q,v\}}$ and $U^{\{q,v\}}$ can be expressed as:
$f=\text{whitening}(\text{Encoder}(q;d))$ (2)
$x=\text{concat}(f,i_{k}^{\{q,v\}}~{}|~{}i_{k}^{q}=2k,i_{k}^{v}=2k+1)$ (3)
$D^{\{q,v\}},U^{\{q,v\}}=g_{D}(x),g_{U}(x)$ (4)
where the Encoder represents the encoder of the model, whitening is a dimensionality reduction algorithm, and concat means to splice the content. $g_{D}$ and $g_{U}$ denote the descending and ascending dimensions of hypernetwork. More formally,
$g(x)=\text{MM}(\text{ReLU}(\text{MM},W_{d}),W_{u})$ (5)
where MM stands for matrix multiplication, ReLU is a activation function.

### v4

# Awakening Augmented Generation: Learning to Awaken Internal Knowledge of Large Language Models for Question Answering

@@ Abstract
Retrieval-Augmented-Generation and Genera-tion-Augmented-Generation have been proposed to enhance the knowledge required for question answering with Large Language Models (LLMs) by leveraging richer context. However, the former relies on external resources, and both require incorporating explicit documents into the context, which increases execution costs and susceptibility to noise data during inference. Recent works indicate that LLMs model rich knowledge, but it is often not effectively activated and awakened. Inspired by this, we propose a novel knowledge-augmented framework, Awakening-Augmented-Generation (AAG), which mimics the human ability to answer questions using only thinking and recalling to compensate for knowledge gaps, thereby awaking relevant knowledge in LLMs without relying on external resources. AAG consists of two key components for awakening richer context. Explicit awakening fine-tunes a context generator to create a synthetic, compressed document that functions as symbolic context. Implicit awakening utilizes a hypernetwork to generate adapters based on the question and synthetic document, which are inserted into LLMs to serve as parameter context. Experimental results on three datasets demonstrate that AAG exhibits significant advantages in both open-domain and closed-book settings, as well as in out-of-distribution generalization. Our code will be available at https://github.com/Xnhyacinth/IAG.
Awakening Augmented Generation: Learning to Awaken Internal Knowledge of Large Language Models for Question Answering
Huanxuan Liao1,2, Shizhu He1,2, Yao Xu1,2, Yuanzhe Zhang1,2, Kang Liu1,2, Shengping Liu3, Jun Zhao1,2 1 The Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences, Beijing, China 2 School of Artificial Intelligence, University of Chinese Academy of Sciences, Beijing, China 3 Unisound, Beijing, China liaohuanxuan2023@ia.ac.cn {yao.xu, shizhu.he, kliu, jzhao}@nlpr.ia.ac.cn
@@ 1 Introduction
We can know more than we can tell. — Michael Polanyi
@@
Figure 1: Compared with RAG and GAG, the proposed AAG eschews external resources, generates a dummy document (explicit awakening) and creates flexible adapters (implicit awakening) for each question.
@@
In recent years, to address hallucinations in LLMs and enhance performance in question answering, researchers have developed several knowledge-augmented methods for LLMs. These methods primarily fall into two categories: Retrieval-Augmented Generation (RAG) Guu et al. (2020) which retrieves documents from external resources (e.g., Wikipedia) and incorporates both the retrieved documents and the question into LLMs Izacard and Grave (2021) (top part of Figure 1). Generation-Augmented Generation (GAG) Kim et al. (2024) which utilizes LLMs such as ChatGPT Ouyang et al. (2022) to generate more relevant documents, which are then used to enhance the answer generation (middle part of Figure 1).
@@
However, these methods have the following disadvantages [^1] A more intuitive comparison can be seen in A.1.: 1) Dependence on external resources, RAG relies on external domain knowledge resources Ke et al. (2024), while GAG depends on a more powerful external LLM as a knowledge generator. This reliance limits their broader application. 2) Increased execution costs, the computing resources and inference time required increase significantly with the number of documents. For example, the typical RAG method FiD Izacard and Grave (2021) must handle over 12,000 tokens to retrieve 100 documents, resulting in more than a $100$-fold increase in prompt length and over $100^{2}$-fold increase in inference time Liu et al. (2023a). Similarly, the GAG method Yu et al. (2023) incurs additional financial costs, such as API calls. 3) Specific retraining, these approaches often require retraining for different domains, tasks and datasets Li et al. (2024). This heightens the challenge of reusing models across different scenarios, resulting in resource inefficiency due to low parameter effectiveness and the need for extensive data.
@@
In fact, LLMs inherently possess rich knowledge and significant potential for tackling knowledge-intensive tasks Bhagavatula et al. (2020). Performance on specific tasks can be improved by more effectively activating and awakening relevant knowledge without external resources. For instance, strategies such as repeating the question twice Xu et al. (2023), consolidating knowledge with prompts like "As far as I know" Yao et al. (2023), and employing visual-language models to imagine images Tang et al. (2023) can all enhance the performance of LLMs on downstream tasks. That is, LLMs model rich knowledge, but it is often not effectively activated and awakened.
Inspired by the above findings and to alleviate the challenges in RAG and GAG, we propose a novel knowledge-augmented framework called Awakening-Augmented Generation (AAG) which emulates the human ability to compensate for knowledge deficits through thinking and recalling in QA. AAG utilizes the context generator to generate a compressed dummy document as symbolic context while reducing computational demands. For instance, AAG uses "official language … Jamaica" (just 20 tokens) as knowledge instead of "Jamaica is regarded… official language is English…" (>200 tokens) in RAG or GAG for the question "what does jamaican people speak?" in WebQ Berant et al. (2013). Additionally, AAG uses the hypernetwork to generate adapters as parameter context for each question, which integrates the advantages of instruction-based learning with parameter-efficient modules to awaken a richer context in LLMs (bottom part of Figure 1).
Specifically, to sufficiently awaken the inherent knowledge of LLMs, we design two main modules to obtain different types of contexts and improve the utilization of relevant knowledge in LLM. The explicit awakening module first employs symbol distillation to compress context, followed by fine-tuning the context generator to generate a concise dummy document, effectively reducing the length of text processing. Next, within the knowledge distillation framework, the implicit awakening module utilizes a hypernetwork to convert questions and other task data (e.g., documents) into adapters inserted into LLMs. This dynamic generation allows for more adaptable and contextually relevant module generation, enhancing the model’s ability to handle diverse and complex tasks effectively. The core idea of AAG is to enable student models that lack rich contextual information to mimic teacher models that possess such information.
@@
We evaluate the proposed AAG on various LLMs, including T5 Roberts et al. (2020a) and Llama2 Touvron et al. (2023). The experimental results across NQ Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017) and WebQ datasets indicate that the proposed AAG yields performance gains while reducing computational expenses and time during inference. Notably, it outperforms baselines that retrieve and generate knowledge $2\%$ under the same document settings and can achieve similar performance while reducing inference cost (tokens processed) by up to 4×4$\times$4 ×. In conclusion, the contributions of this paper are summarized as follows:
@@
  • We propose a new knowledge augmentation framework AAG to awaken richer context (symbolic and parameter context) more efficiently without relying on external resources.
  • We make use of a text-conditioned hypernetwork to generate parameter-efficient modules as parameter context based on the question and a dummy compressed document.
@@
  • Experimental results indicate that AAG effectively awakens the relevant knowledge of LLMs which demonstrates significant advantages in both open-domain and closed-book settings while reducing inference cost.
@@ 2 Related Work
Knowledge Enhancement has usually been adopted to alleviate the issue of insufficient knowledge in LLMs. There are two main methods: RAG Sun et al. (2019); Wang et al. (2024) and GAG Abdallah and Jatowt (2023). The typical RAG method FiD Izacard and Grave (2021) retrieves documents from Wikipedia to answer questions. LLMs serving as a knowledge base have been the focus of numerous studies that advocate the extraction of knowledge from such models (e.g., GPT-3). For instance, Yu et al. (2023) generates 10 documents for each question. However, RAG requires external resources, and both RAG and GAG need verbose long contexts. Recently, methods have been developed to enhance LLMs’ abilities by simulating human imagination of visual information using existing visual-language models Tang et al. (2023); Akter et al. (2024). Our proposed method not only eliminates the need for external resources but also improves the efficiency of activating internal knowledge within LLMs.
@@
Knowledge Distillation is a technique where a smaller model learns to mimic the predictions of a larger model, aiming to retain performance while reducing computational resources Hinton et al. (2015). Recent studies West et al. (2022) present symbolic knowledge distillation, a process that facilitates knowledge transfer from a teacher model via extracting training data to subsequently train a student model Wang et al. (2023b); Ranaldi and Freitas (2024). In this paper, the process of obtaining compressed context during context generator fine-tuning resembles a form of symbolic distillation. Regarding training, our emphasis lies in distilling the long-context modeling abilities of LLMs.
Hypernetworks is designed to reduce the number of parameters Ha et al. (2016), i.e., a small neural network generates parameters for another big neural network. It offers a solution that reduces the dependency on gradient descent for specific domains. Recent studies Phang et al. (2022); Ivison et al. (2023) have explored the enhancement of model performance in zero/few-shot settings through meta-learning involving hypernetworks. We utilize hypernetworks to acquire parameter context by dynamically converting the question and the other data to adapters inserted into LLMs for efficiency and generalization.
@@ 3 Method
In this section, we introduce the details of AAG to activate LLMs’ intrinsic knowledge and obtain a richer context for QA. The fundamental premise underlying this method is that QA with a richer context (teacher model) yields a better internal representation and greater performance (e.g., RAG with retrieved documents). Therefore, to enable a student model without external documents as context to also possess rich context, it is necessary to both learn to independently generate context (though not excessively long) and to allow the student model to mimic and acquire rich internal representations.
Specifically, as shown in Figure 2, AAG comprises two main modules. Explicit awakening with long context compression learns to generate a compressed dummy document (§ 3.2). Implicit awakening with the hypernetwork leverages hidden knowledge that learns a shared knowledge feature projection across questions (§ 3.3). The hypernetwork is trained to generate lightweight LoRA modules to align the question and the internal knowledge. Besides, there is long context distillation in training, which learns the teacher’s rich representations to compensate for missing knowledge in label learning (§ 3.4).
Figure 2: Overview of AAG method. In the inference phase, for each question, the explicit awakening (context generator) generates a short dummy document and the implicit awakening (hypernetwork) generates a specific LoRA module. During training, there are two stages: the first stage is the pre-training of the context generator (§ 3.2), aiming at its ability to imagine a short dummy document based on the question, and the second stage is the hypernetwork fine-tuning (§ 3.3) using long context distillation (§ 3.4) to obtain a question-specific LoRA module.
@@ 3.1 Formulation
The formulation of our task follows RAG for QA Guu et al. (2020). Let $\mathcal{V}^{*}$ denote the infinite set, encompassing all potential strings over the tokens in vocabulary $\mathcal{V}$, and this includes the empty string. An instance within a QA dataset is defined as a triplet $(\boldsymbol{q},\boldsymbol{a},\boldsymbol{c})$ comprising question $\boldsymbol{q}$, answer $\boldsymbol{a}$, and context $c$, where $\boldsymbol{q},\boldsymbol{a},\boldsymbol{c}\in\mathcal{V}^{*}$. Conventionally, the context $c$ is drawn from the knowledge corpus $\mathcal{Z}$, like Wikipedia, whereby $\mathcal{Z}\subset\mathcal{V}^{*}$. Additional background details are available in B.1.
3.2 Explicit Awakening with Context Generator
To obtain the short dummy document $\boldsymbol{d}$, we fine-tune a context generator  [^2] We discuss the role of context generator in the A.2. to utilize its knowledge in generating a compressed dummy document as symbolic context, thereby reducing input length. Simultaneously, we avoid dependence on a fixed knowledge base and minimize knowledge corpus errors by incorporating potentially useful context Lee et al. (2023). Employing a knowledge distillation framework, the student model learns to generate the compressed text that the teacher model produces based on extensive context.
Specifically, for each data point $\mathcal{D}_{\text{train}}=\{(\boldsymbol{q}_{i},\boldsymbol{a}_{i},\boldsymbol{c}_{i})\}_{i=1}^{n}$, we apply the long-context compression method LongLLMLingua Jiang et al. (2023) to the retrieved text $\boldsymbol{c}_{i}$, resulting in the compressed text $\boldsymbol{c^{\prime}}_{i}$. As shown in the left part of Figure 2, subsequently, we fine-tune the context generator $p_{\theta}$ with trainable parameters $\theta$ to fully leverage its inherent knowledge for generating $\boldsymbol{c^{\prime}}_{i}$, which guides the model to think about its knowledge and generate a short dummy document. Our objective is to minimize the negative log-likelihood of the compressed text $\boldsymbol{c^{\prime}}_{i}$ sequence given the specific prompt $\boldsymbol{p}$ (B.2) and the question $\boldsymbol{q}_{i}$.
$\mathcal{L}_{ce}=-\frac{1}{n}\sum_{i=1}^{n}\log p_{\theta}(\boldsymbol{c^{\prime}}_{i}\mid\boldsymbol{p},\boldsymbol{q}_{i})$ (1)
This process enables LLMs to conceive compressed document that robustly parallels the question’s knowledge requirements.
3.3 Implicit Awakening with Hypernetwork
@@
We utilize the hypernetwork [^3] We conduct a detailed analysis of the reasons behind the hypernetwork in the A.3. to convert the question $\boldsymbol{q}$ and short dummy document $\boldsymbol{d}$ into a specific parameter-efficient LoRA module inserted into the LLM, serving as the parameter context for the question. This is akin to repeating the question in the prompt Xu et al. (2023) and incorporating certain topical cues to stimulate the model’s recall of relevant questions Wang et al. (2023c). However, the distinction lies in the fact that they serve as wake-up features, whereas we are generating model parameters as knowledge awakening.
The hypernetwork architecture for generating LoRA weights is detailed in Figure 3. Specifically, $D_{k}^{q}$ and $U_{k}^{q}$ represent the low-rank projections of layer $k$ associated with the $\mathcal{Q}$, while $D_{k}^{v}$ and $U_{k}^{v}$ correspond to those associated with the $\mathcal{V}$. The hypernetwork, denoted as $g_{D}$ and $g_{U}$, takes $\textit{concat}(f,i_{k}^{\{q,v\}})$ as input, where $f$ is the feature vector obtained using the model’s encoder and reduced in dimensionality via a whitening algorithm Su (2021). The term $idx_{k}^{\{q,v\}}\in\{0,\ldots,2\times\#\text{blocks}\}$ signifies the positional embedding, differentiating between layers and $\mathcal{QV}$. Each hypernetwork is characterized by weights $W_{d}$ and $W_{u}$, representing the down and up projections, respectively. The hypernetwork equations for $D^{\{q,v\}}$ is expressed as follows:
$f_{i}=\text{whitening}(\text{Encoder}(\boldsymbol{q}_{i};\boldsymbol{d}_{i}))$ (2)
$g(x)=W_{u}\cdot\text{ReLU}(W_{d}\cdot x)$ (3)
$D^{\{q,v\}}=g_{D}((f_{i};idx_{k}^{\{q,v\}}))$ (4)
where Encoder represents the encoder of the model, whitening is a dimensionality reduction algorithm, ReLU is an activation function, and $idx_{k}^{q}=2k,idx_{k}^{v}=2k+1$. $g_{D}$ and $g_{U}$ represent the dimension reduction and dimension increase functions of the hypernetwork, respectively.